# Multimodal 3D Genome Pre-training

### v1

@@ 5.3 Robust Performance Across Cell Types
The utilization of cell-type specific data, i.e., Hi-C contact maps and epigenomic tracks, empowers MIX-HIC to achieve accurate predictions for novel cell types. We perform a cross-dataset evaluation with the GM12878 and K562 cell lines on two downstream tasks, including the prediction of Hi-C contact maps and chromatin loops, to assess the generalization ability of MIX-HIC. In particular, all the models are trained on one cell line and evaluated on the other. The evaluation results of Hi-C contact map prediction and chromatin loops prediction tasks are shown in Figure 5. We note a decrease in performance when all models are evaluated on external datasets compared to within-dataset testing (see Tables 2 and 3). Nevertheless, MIX-HIC continues to outperform other methods, indicating its strong generalization ability in real-world scenarios.
@@ 5.4 Ablation Study
The MIX-HIC consists of three fundamental components: the integration of Hi-C contact maps with epigenomic tracks, the proposed cross-modal interaction and cross-modal mapping blocks, and the self-supervised pre-training strategy. We conduct an ablation study to assess their effectiveness. Specifically, we investigate: (1) bimodal versus unimodal representations, (2) the effectiveness of missing modality features complement, and (3) the impact of pre-training versus the absence of pre-training.
The comparison results for three tasks on two datasets are displayed in Figure 6. First, we note that in the chromatin loop prediction task on the K562 dataset, the naive fusion of bimodal representations without loss constraints, denoted as MIX-HIC-Bimodal w/o pre-train, results in performance degradation compared to using the Hi-C contact map alone (referred to as Map only). This may be attributed to the inherent data heterogeneity between the two modalities. With self-supervised pre-training, the integration of both epigenomic tracks and Hi-C contact maps (labeled as MIX-HIC-Bimodal) consistently outperforms the unimodal approach, providing unified and comprehensive representations. Second, MIX-HIC effectively captures the implicit structural connections between bimodal data, demonstrating enhanced performance through the utilization of complementary features (designated as MIX-HIC-InferMap) compared to approaches relying solely on epigenomic tracks (denoted as Epi only). This demonstrates the effectiveness of MIX-HIC in inferring missing features, offering a potential solution to address the data scarcity in 3D genomic studies. Finally, the self-supervised pre-training empowers MIX-HIC to learn a robust and unified representation, achieving superior generalization capability compared to non-pre-trained models.
Figure 6: Ablation study results of MIX-HIC for three downstream tasks.
@@ 6 Conclusion
In this work, we present MIX-HIC, a novel multimodal foundation model for diverse 3D genome downstream tasks by integrating Hi-C contact maps and epigenomic tracks. To facilitate a unified and comprehensive understanding of 3D genome organization, we constructed the largest paired 3D genome dataset to date, comprising over one million high-quality samples. As the first multimodal foundation model in this domain, MIX-HIC incorporates two key innovations: (1) a cross-modal interaction block that jointly learns modal-invariant and modal-specific representations, effectively capturing both shared and unique biological patterns across modalities; and (2) a cross-modal mapping block that regularizes the multimodal feature space and enables robust imputation of missing modalities features, alleviating the practical challenges posed by the high costs of Hi-C contact map acquisition. Comprehensive experimental results demonstrate that MIX-HIC achieves state-of-the-art performance on three key downstream tasks across two cell lines. We anticipate this work establishes a new paradigm for multimodal 3D genomics analysis and paves the way for future advancements in the field.
@@
In this section, we present the implementation details of MIX-HIC, including data source and processing, whole-chromosome loop annotation, as well as experimental settings. Note that we will open-source our code and dataset after the paper publishing.
Table S1: Data sources and accession numbers from 4DN Data Portal and ENCODE Data Portal.
@@ A.1 Data Source and Processing
Detailed reference numbers for all data sources are provided in Table S1. We download the BAM files for epigenomic tracks and convert them into bigWig files using deepTools (Ramírez et al., 2016) with RPGC normalization. Two 250 kb epigenomic regions, corresponding to the x-axis and y-axis of a Hi-C contact map, are first averaged over every 100 bp and transformed using $log(x+1)$ to reduce data variability, and then the processed regions are concatenated to form the final epigenomic tracks input with a length of 5,000. The Hi-C contact maps are binned at 5 kb resolution and normalized using KR normalization, and then divided into $50\times 50$ sub-matrices based on the loci of each sample.
The training samples of three downstream tasks are generated using two widely adopted cell lines, GM12878 and K562, as described below. First, the Hi-C contact maps and epigenomic tracks within a 250 kb genomic region upstream and downstream of the gene transcription start site (TSS) are used to predict CAGE-seq expression, with TSS annotations obtained from previous work (Karbalayghareh et al., 2022). For this task, the number of training samples for GM12878 and K562 are 16,046 and 14,739, respectively. Second, the same pairwise samples of Hi-C contact maps and epigenomic tracks are used for the Hi-C contact map prediction task, with the same sample sizes as mentioned above. Third, the positive chromatin loops (1,344,270 for GM12878 and 137,558 for K562) are derived from the ChIA-PET data, while an equal number of negative loops are randomly sampled based on two criteria following prior research (Wang et al., 2022): (1) matching the distance distribution of positive samples using the distances probability density function, and (2) selecting interactions with distances greater than the maximum distance observed in the positive samples to enhance the diversity of the negative samples. Following previous work (Wang et al., 2022; Tan et al., 2023; Yang et al., 2023), we partition the chromosomes into distinct training, validation, and test sets. Specifically, chromosomes 10 and 11 serve as the validation set, chromosomes 3, 13, and 17 as the test set, and the remaining chromosomes are used for model training across three downstream tasks.
@@ A.2 Whole-chromosome Loops Annotation
The well-trained MIX-HIC model is capable of predicting all potential chromatin interactions across individual chromosomes. The detection of chromatin loops within each chromosome consists of two key steps: (1) scoring potential chromatin interactions using MIX-HIC and (2) aggregating loops through a clustering algorithm. First, candidate elements are identified by examining the diagonals of the raw Hi-C contact matrix, where observed interaction frequencies are statistically compared to expected values based on a Poisson distribution, retaining only elements with p-values below 0.01. Subsequently, the trained MIX-HIC model is employed to predict loop probability scores, with scores exceeding 0.5 referred to as candidate loops. These candidates are further grouped to identify significant loops using a density-based clustering algorithm following (Zhang & Blanchette, 2022). For each candidate, a local density score and a minimum distance metric are computed. Candidates exhibiting low distance values are subsequently eliminated to minimize redundancy. Cluster centers are determined via a target-decoy search, ensuring that the false discovery rate remains below 5%. Within each cluster, the candidate demonstrating the highest local density is designated as the representative loop.
@@ A.3 Experimental Settings
During the pre-training stage, MIX-HIC is configured with 500 epochs, a learning rate of 1e-5, and a batch size of 256. For the CAGE-seq expression prediction task, the predefined feature dimension $C$ and learning rate are set to 256 and 1e-4, respectively. For the prediction of Hi-C contact maps and chromatin loops, these parameters are specified as 128 and 1e-5, respectively. The number of transformer blocks $T$ in each encoder, contact map-grounded fusion block, and decoder is set to 2. Fine-tuning is conducted with a batch size of 64, utilizing the AdamW optimizer (Loshchilov & Hutter, 2017) with the momentum parameters $\beta_{1}$ and $\beta_{2}$ initialized to 0.9 and 0.999, respectively. The fine-tuning process is configured with a maximum of 200 epochs, and an early stopping strategy is employed with a patience parameter of 20 to prevent overfitting.
@@
Simple alignment of multimodal features can result in information loss, which may compromise the performance of downstream tasks. Under conditions of perfect feature alignment, the prediction error using aligned features is at least $\Gamma_{q}$ greater than that using the raw inputs. We generalize the theorem from (Dong et al., 2023) to the case of two modalities as follows.
Let $q$ represent the joint distribution of $(z_{1},z_{2},t)$, where $z_{1}$ and $z_{2}$ denote the two modalities, and $t$ is the target variable. The information gap is formulated as $\Gamma_{q}:=\max\{U(z_{1};t),U(z_{2};t)\}-\min\{U(z_{1};t),U(z_{2};t)\}$, with $U(z_{j};t)$ representing the mutual information between the target $t$ and the modality $z_{j}$. The methodology utilizes the cross entropy loss $\mathcal{L}_{\text{CE}}$, two feature encoders $f^{1}(\cdot)$ and $f^{2}(\cdot)$, a feature prediction function $h(\cdot)$, and a raw data prediction function $h^{\prime}(\cdot)$. If the bimodal features $\mathbf{F}^{1}=f^{1}(z_{1})$ and $\mathbf{F}^{2}=f^{2}(z_{2})$ are perfectly aligned, i.e., $\mathbf{F}^{1}=\mathbf{F}^{2}$, then we can obtain the theorem:
This theorem implies that if one modality is more informative than the others (i.e., there exists a significant information gap), perfect alignment may lead to a substantial increase in prediction error. The underlying reason is that perfect alignment constrains the aligned features to preserve only the predictive information shared across all modalities, which may result in the loss of modal-specific information that is potentially critical for achieving accurate predictions. The proof of this theorem is presented as follows:
@@
$U(\mathbf{F}^{1},\mathbf{F}^{2};t)=U(\mathbf{F}^{1};t)+U(\mathbf{F}^{2};t|\mathbf{F}^{1})=U(\mathbf{F}^{2};t)+U(\mathbf{F}^{1};t|\mathbf{F}^{2}).$ (10)
@@
By applying the data processing inequality (Beaudry & Renner, 2012): $U(\mathbf{F}^{1};t)\leq U(z_{1};t)$ and $U(\mathbf{F}^{2};t)\leq U(z_{2};t)$, we derive:
$U(\mathbf{F}^{1},\mathbf{F}^{2};t)=\min\{U({F}^{1};t),U({F}^{2};t)\}\leq\min\{U(z_{1};t),U(z_{2};t)\}\leq\max\{U(z_{1};t),U(z_{2};t)\}\leq U(z_{1},z_{2};t).$ (12)
According to the variational form of conditional entropy $H(t|z_{1},z_{2})=\inf_{h}\mathbb{E}_{q}[\mathcal{L}_{\text{CE}}(h(z_{1},z_{2}),t)]$ (Zhao et al., 2022) and the definition of mutual information $U(X;Y)=H(Y)-H(X|Y)$, we can conclude the theorem as follows:
@@ C.1 Proportion of Validated Loops versus Predicted Loops
Table S2 presents a comparative analysis of chromatin interaction loops identified through ChIA-PET experiments and those predicted by different approaches.
Table S2: Comparison of ChIA-PET-supported loops with those identified by various deep learning methods.
@@
ChromoSight 3845 6044 0.64 1247 1963 0.63
Peakachu 1992 2904 0.69 718 1221 0.59
DLoopCaller 3849 6019 0.64 1048 2274 0.46
MIX-HIC 5179 5893 0.88 1064 1588 0.67
@@ C.2 Example of Whole-Chromosome Chromatin Loop Comparison
Figure S1 illustrates the performance evaluation of chromatin loops predicted by various methods (black points, lower left) against experimentally validated interactions (red points, upper right). Our predictions (the first left panel of Figure S1) demonstrate significantly fewer false positive loops compared to experimentally validated ChIA-PET interactions in the Chromosome 13 region (20Mb-22Mb) from the GM12878 dataset, as evidenced by the correspondent counts between red and black points. The comparative visualization highlights the enhanced specificity of our approach in identifying chromatin interactions across the genomic region of interest.
[Uncaptioned image] Figure S1: Comparison of predicted chromatin loops (black points, lower left) and experimentally validated ChIA-PET interactions (red points, upper right) in the Chromosome 13 region (20Mb-22Mb) from the GM12878 dataset.
@@ C.3 Hyperparameter Analysis
The number of Transformer $T$ blocks in each layer of the encoder, contact map-grounded fusion block, and decoder, along with the predefined feature dimension $C$ play a pivotal role in MIX-HIC. We evaluate the impact of this parameter by testing $C=\{64,128,256\}$ and $T=\{2,4,8\}$ as depicted in Figure S2 and Figure S3, respectively. The highest performance is achieved at $C=128$ for HI-C contact maps and chromatin loops prediction tasks, while $C=256$ performs best for CAGE-seq expression prediction. Additionally, $T=2$ demonstrates the most robust performance across all three downstream tasks. MIX-HIC exhibits strong resilience to parameter variations, underscoring its robustness. These experimental results substantiate the parameter configurations employed in MIX-HIC.
@@
[Uncaptioned image] Figure S2: The performance across different predefined feature dimensions $C=\{64,128,256\}$ for three downstream tasks in GM12878 and K562 cell lines.

### v2

@@ 5.3 Robust Performance Across Cell Types
The utilization of cell-type specific data, i.e., Hi-C contact maps and epigenomic tracks, empowers MIX-HIC to achieve accurate predictions for novel cell types. We perform a cross-dataset evaluation with the GM12878 and K562 cell lines on the chromatin loop detection task to assess the generalization ability of MIX-HIC. In particular, all the models are trained on one cell line and evaluated on the other. The evaluation results of cross-cell type prediction are shown in Figure 5. We note a decrease in performance when all models are evaluated on external datasets compared to within-dataset testing (see Table 3). Nevertheless, MIX-HIC-Bimodal continues to outperform other methods, indicating its strong generalization ability in real-world scenarios.
Table 5: AUROC results of ablation studies on loss terms for chromatin loop detection task. Symbols ‘✓’ and ‘-’ denote present and absent, respectively.
$\mathcal{L}_{\text{con}}$ $\mathcal{L}_{\text{orth}}$ $\mathcal{L}_{\text{mapping}}$ GM12878 K562
- - 0.9136 0.9099
- 0.9183 0.9156
0.9209 0.9194
@@ 5.4 Ablation Study
To evaluate the effectiveness of key components in MIX-HIC, we perform ablation studies focusing on the proposed loss terms and the representation learning ability across multimodal data.
Three critical loss terms are employed during MIX-HIC pre-training (i.e., $\mathcal{L}_{\text{con}}$, $\mathcal{L}_{\text{orth}}$, and $\mathcal{L}_{\text{mapping}}$). We examine the effects of each loss component. As shown in Table 5, the model equipped with $\mathcal{L}_{\text{con}}$ provides a strong baseline. The orthogonal loss enhances discrimination between modal-invariant and modal-specific representations, contributing about 0.5% AUROC improvement compared to simply feature alignment, as demonstrated in Theorem 1. Further details for the orthogonal constaint are provided in Appendix C.4. Although the cross-modal mapping loss provides modest enhancement, it enables missing modality inference of MIX-HIC. These findings demonstrate that robust multimodal fusion is achieved by explicitly separating modal-invariant and modal-specific representations using appropriate loss terms.
Table 6: Ablation results of each modality, with values reported as $R^{2}$ (Hi-C contact map prediction), AUROC (chromatin loops dectection), and $R^{2}$ (CAGE-seq expression prediction). Symbol $\circ$’ represents inferred embeddings from the other modality.
Tasks Epi. Hi-C Pre-trained GM12878 K562
- - 0.8481 0.7709
Hi-C contact map prediction $\circ$ 0.8724 0.8001
- - 0.8236 0.8054
$\circ$ 0.8494 0.8226
- - 0.9065 0.9072
$\circ$ 0.9135 0.9159
- 0.9091 0.8859
Chromatin loops dectection 0.9209 0.9194
- - 0.8514 0.8710
$\circ$ 0.8684 0.8870
- 0.8614 0.8755
CAGE-seq expression prediction 0.8833 0.9077
In addition, we assess the representation learning ability of MIX-HIC across three aspects: the contribution of pre-training, the superiority of multimodal over unimodal representations, and the efficiency of modality completion. For the Hi-C contact map prediction task, the model is trained to predict the contact map using only 1D epigenomic tracks as input. The Hi-C contact map itself is the prediction target (output), making its use as an input feature methodologically invalid. Detailed comparison results appear in Table 6. The pre-trained bimodal MIX-HIC achieves superior performance compared to its non-pre-trained counterpart, confirming pre-training’s benefit for unified representations. Moreover, while pre-trained bimodal MIX-HIC consistently outperforms single-modal variants, the non-pre-trained bimodal version underperforms compared to using Hi-C data alone for K562 chromatin loop detection due to heterogeneity between bimodal data, highlighting the importance of learning both modality-specific information and general patterns. Finally, when Hi-C data is unavailable, combining epigenomic embeddings with inferred Hi-C embeddings improves performance over using epigenomic data alone, validating the cross-modal mapping block’s effectiveness for data scarcity scenarios.
@@ 6 Conclusion
In this work, we present MIX-HIC, a novel multimodal foundation model for diverse 3D genome downstream tasks by integrating Hi-C contact maps and epigenomic tracks. To facilitate a unified and comprehensive understanding of 3D genome organization, we construct the largest paired 3D genome dataset to date, comprising over one million high-quality samples. As the first multimodal foundation model in this domain, MIX-HIC incorporates two key innovations: (1) a cross-modal interaction block that jointly learns modal-invariant and modal-specific representations, effectively capturing both shared and unique biological patterns across modalities; and (2) a cross-modal mapping block that regularizes the multimodal feature space and enables robust imputation of missing modalities features, alleviating the practical challenges posed by the high costs of Hi-C contact map acquisition. Comprehensive experimental results demonstrate that MIX-HIC achieves state-of-the-art performance on three key downstream tasks across two cell lines.
7 Acknowledgments
This work was supported in part by the National Key R&D Program of China (Grant No.2023YFF0725001), in part by the National Natural Science Foundation of China (Grant Nos.92370204, 62572413), in part by the guangdong Basic and Applied Basic Research Foundation (Grant Nos.2023B1515120057, 2025A1515012944), and in part by the Education Bureau of Guangzhou.
@@ 7 Acknowledgments
In this section, we present the implementation details of MIX-HIC, including data source and processing, whole-chromosome loop annotation, as well as experimental settings.
Table 7: Data sources and accession numbers from 4DN Data Portal and ENCODE Data Portal.
@@ A.1 Data Source and Processing
Detailed reference numbers for all data sources are provided in Table 7. We download the BAM files for epigenomic tracks and convert them into bigWig files using deepTools [40] with RPGC normalization. Two 250 kb epigenomic regions, corresponding to the x-axis and y-axis of a Hi-C contact map, are first averaged over every 100 bp and transformed using $log(x+1)$ to reduce data variability, and then the processed regions are concatenated to form the final epigenomic tracks input with a length of 5,000. The Hi-C contact maps are binned at 5 kb resolution and normalized using KR normalization, and then divided into $50\times 50$ sub-matrices based on the loci of each sample.
As the first attempt to create the multimodal 3D genome foundation model, we prioritize epigenomic tracks with ATAC-seq and DNase-seq because they capture the genome’s foundational regulatory information. While DNase-seq and ATAC-seq have some inherent redundancy, they also offer complementary insights due to their distinct enzymatic biases. The enzymes’ different cutting biases cause each assay to detect unique accessible sites, which together yield a more complete accessibility landscape [43].
The training samples of three downstream tasks are generated using two widely adopted cell lines, GM12878 and K562, as described below. First, the Hi-C contact maps and epigenomic tracks within a 250 kb genomic region upstream and downstream of the gene transcription start site (TSS) are used to predict CAGE-seq expression, with TSS annotations obtained from previous work [12]. For this task, the number of training samples for GM12878 and K562 are 16,046 and 14,739, respectively. Second, the same pairwise samples of Hi-C contact maps and epigenomic tracks are used for the Hi-C contact map prediction task, with the same sample sizes as mentioned above. Third, the positive chromatin loops (1,344,270 for GM12878 and 137,558 for K562) are derived from the ChIA-PET data, while an equal number of negative loops are randomly sampled based on two criteria following prior research [10]: (1) matching the distance distribution of positive samples using the distances probability density function, and (2) selecting interactions with distances greater than the maximum distance observed in the positive samples to enhance the diversity of the negative samples. Following previous work [10, 6, 5], we partition the chromosomes into distinct training, validation, and test sets. Specifically, chromosomes 10 and 11 serve as the validation set, chromosomes 3, 13, and 17 as the test set, and the remaining chromosomes are used for model training across three downstream tasks.
@@ A.2 Whole-chromosome Loops Annotation
The well-trained MIX-HIC model is capable of predicting all potential chromatin interactions across individual chromosomes. The detection of chromatin loops within each chromosome consists of two key steps: (1) scoring potential chromatin interactions using MIX-HIC and (2) aggregating loops through a clustering algorithm. First, candidate elements are identified by examining the diagonals of the raw Hi-C contact matrix, where observed interaction frequencies are statistically compared to expected values based on a Poisson distribution, retaining only elements with p-values below 0.01. Subsequently, the trained MIX-HIC model is employed to predict loop probability scores, with scores exceeding 0.5 referred to as candidate loops. These candidates are further grouped to identify significant loops using a density-based clustering algorithm following [44]. For each candidate, a local density score and a minimum distance metric are computed. Candidates exhibiting low distance values are subsequently eliminated to minimize redundancy. Cluster centers are determined via a target-decoy search, ensuring that the false discovery rate remains below 5%. Within each cluster, the candidate demonstrating the highest local density is designated as the representative loop.
@@ A.3 Experimental Settings
During the pre-training stage, MIX-HIC is configured with 500 epochs, a learning rate of 1e-5, and a batch size of 256. For the CAGE-seq expression prediction task, the predefined feature dimension $C$ and learning rate are set to 256 and 1e-4, respectively. For the prediction of Hi-C contact maps and chromatin loops, these parameters are specified as 128 and 1e-5, respectively. The number of transformer blocks $T$ in each encoder, contact map-grounded fusion block, and decoder is set to 2. Fine-tuning is conducted with a batch size of 64, utilizing the AdamW optimizer [45] with the momentum parameters $\beta_{1}$ and $\beta_{2}$ initialized to 0.9 and 0.999, respectively. The fine-tuning process is configured with a maximum of 200 epochs, and an early stopping strategy is employed with a patience parameter of 20 to prevent overfitting.
@@
Simple alignment of multimodal features can result in information loss, which may compromise the performance of downstream tasks. Under conditions of perfect feature alignment, the prediction error using aligned features is at least $\Gamma_{q}$ greater than that using the raw inputs. Based on [36], we generalize the theorem to the case of two modalities as Theorem 1. This theorem implies that if one modality is more informative than the others (i.e., there exists a significant information gap), perfect alignment may lead to a substantial increase in prediction error. The underlying reason is that perfect alignment constrains the aligned features to preserve only the predictive information shared across all modalities, which may result in the loss of modal-specific information that is potentially critical for achieving accurate predictions. The proof of this theorem is presented as follows:
@@
$U(\mathbf{F}^{1},\mathbf{F}^{2};t)$ $=U(\mathbf{F}^{1};t)+U(\mathbf{F}^{2};t|\mathbf{F}^{1})$ (10)
$=U(\mathbf{F}^{2};t)+U(\mathbf{F}^{1};t|\mathbf{F}^{2}).$ (11)
@@
By applying the data processing inequality [46]:
$U(\mathbf{F}^{1};t)\leq U(z_{1};t);U(\mathbf{F}^{2};t)\leq U(z_{2};t),$ (13)
we derive $U(\mathbf{F}^{1},\mathbf{F}^{2};t)$ as follows:
$U(\mathbf{F}^{1},\mathbf{F}^{2};t)$ $=\min\{U({F}^{1};t),U({F}^{2};t)\}$
$\leq\min\{U(z_{1};t),U(z_{2};t)\}$
$\leq\max\{U(z_{1};t),U(z_{2};t)\}$
$\leq U(z_{1},z_{2};t).$
According to the variational form of conditional entropy $H(t|z_{1},z_{2})=\inf_{h}\mathbb{E}_{q}[\mathcal{L}_{\text{CE}}(h(z_{1},z_{2}),t)]$ [47] and the definition of mutual information U(X;Y)=H(Y)H(X|Y)U(X;Y)=H(Y)-H(X|Y), we can conclude the theorem as follows:
@@ C.1 Proportion of Validated Loops versus Predicted Loops
We define the quantitative metric ‘proportion’ as the ratio of experimentally validated loops to computationally predicted loops [44]. Table 8 presents a comparative analysis of chromatin interaction loops identified through ChIA-PET experiments and those predicted by different approaches. MIX-HIC exceeds runner-up methods by 28% and 9% proportion on GM12878 and K562 datasets, respectively.
Table 8: Comparison of ChIA-PET-supported loops with those identified by various deep learning methods.
@@
ChromoSight 3845 6044 0.64 993 1628 0.61
Peakachu 1992 2904 0.69 795 1492 0.53
DLoopCaller 4301 6878 0.62 1048 2274 0.46
[…]
@@ C.2 Example of Whole-Chromosome Chromatin Loop Comparison
Figure 6 illustrates the performance evaluation of chromatin loops predicted by various methods (black points, lower left) against corresponding experimentally validated interactions (red points, upper right). Specifically, we use the same genome regions, chromosome 13: 20Mb-22Mb from the GM12878 dataset, and identical ground truth data for evaluating all compared methods. Our predictions (the first left panel) demonstrate significantly fewer false positive loops compared to experimentally validated ChIA-PET interactions, as evidenced by the correspondent counts between red and black points. Our method achieves a proportion of 0.57, compared to the runner-up’s 0.32, highlighting MIX-HIC’s superior ability to accurately identify chromatin loops while producing fewer false positives.
[Uncaptioned image] Figure 6: Comparison of predicted chromatin loops (black points, lower left) and corresponding experimentally validated ChIA-PET interactions (red points, upper right) in the chromosome 13 region (20Mb-22Mb) from the GM12878 dataset. Unmatched experimentally validated ChIA-PET interactions are represented by green points in the upper right.
@@ C.3 Hyperparameter Analysis
The number of Transformer $T$ blocks in each layer of the encoder, contact map-grounded fusion block, and decoder, along with the predefined feature dimension $C$ play a pivotal role in MIX-HIC. We evaluate the impact of this parameter by testing $C=\{64,128,256\}$ and $T=\{2,4,8\}$ as depicted in Figure 7 and Figure 8, respectively. The highest performance is achieved at $C=128$ for HI-C contact maps and chromatin loops prediction tasks, while $C=256$ performs best for CAGE-seq expression prediction. Additionally, $T=2$ demonstrates the most robust performance across all three downstream tasks. MIX-HIC exhibits strong resilience to parameter variations, demonstrating our model’s robustness and parameter efficiency. The performance suggests our architecture reaches a “sweet spot” with a moderate parameter count and computational efficiency, sufficient to capture the essential biological patterns without incurring the high risk of overfitting during fine-tuning on smaller, task-specific datasets. As our pre-training corpus expands, we anticipate larger architectures will become beneficial. These experimental results substantiate the parameter configurations employed in MIX-HIC.
@@
[Uncaptioned image] Figure 7: The performance across different predefined feature dimensions $C=\{64,128,256\}$ for three downstream tasks in GM12878 and K562 cell lines.